OpenStudy (anonymous):

What equation represents a line that passes through (2, -1) and has a slope of 4

OpenStudy (anonymous):

y=mx+c is the basic equation of a straight line, where m is the gradient (slope) and c is a constant. Since you already know the gradient (4) and you have been given two points (2, -1) you just need to find c. for the equation. To do this, substitute x and y in the equation from the point you've been given and substitute the gradient to find c. So; -1=4*2 +c Rearrange to find c. -1 = 8 +c -9 = c Therefore the equation of the line is: y=4x - 9